If two dice are thrown, find the probability that the sum is neither 5 nor 9?
Show answerHide answer
Correct answer: (D) 7/9
Explanation
A sum of 5 occurs 4 ways and a sum of 9 occurs 4 ways out of 36, so P(5 or 9) = 8/36 = 2/9. The required probability is 1 − 2/9 = 7/9.